Two people departed simultaneously in opposite directions. Later that day they were 18 km apart. If the faster hiker traveled 3 km more than the slower person, how far did the slower person travel?
Answer by oberobic(2304)

distance they are apart = 18 km
.
x = distance walked by the slower hiker
x + 3 = distance walked by the faster hiker
.
x + (x+3) = 18
.
2x +3 = 18
2x = 15
x = 7 1/2
.
x+3 = 10 1/2
.
So, the answer is:
The slower hiker traveled 7 1/2 km, and the faster hiker traveled 10 1/2 km.
.
Done.